OpenStudy (anonymous):

Solve the equation. -5y -9 = -(y - 1)

OpenStudy (anonymous):

here http://akk.li/pics/anne.jpg

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Haha, you're funny (:

OpenStudy (anonymous):

lol

OpenStudy (itsbribro):

-5y-9=-(y-1) = -5/2 = -2.5 y = -2.5